Bananagirl:

Part A Mark the focus of the parabola you are going to create at F(6, 4). Draw a horizontal line that is 6 units below the focus. This line will be the directrix of your parabola. What is the equation of the line?

surjithayer:

4-6=-2 eq. of line (directrix) is y=-2

surjithayer:



surjithayer:

let P(x,y) be any point on the parabola. Draw \[PM~ perpendicular~ on~ directrix ~y=2 ~or~ y-2=0\] then by def. of parabola PF=PM \[\sqrt{(x-4)^2+(y-6)^2}=\frac{ |y-2| }{ \sqrt{0^2+1^2} }\] squaring \[(x-4)^2+(y-6)^2=(y-2)2\] simplify and you get the eq. of parabola.